Solve the system by addition
2x-4y=7
4x+2y=9

Answer by venugopalramana(3286)

2x-4y=7......................I
4x+2y=9.................II
EQN.I+EQN.II*2
2X-4Y+8X+4Y=7+18=25
10X=25
X=25/10=2.5
4*2.5+2Y=9
2Y=9-10=-1
Y=-1/2=-0.5